OpenStudy (yash2651995):

integrate (from 1 to e^(37)) [pi.sin(pi.ln x)]/x

OpenStudy (anonymous):

let \[t=\pi \ln x\]

OpenStudy (yash2651995):

\[\int\limits_{1}^{e ^{37}} [\pi \sin (\pi \ln x)] \div x\]

OpenStudy (anonymous):

try to sub it and tell me what u get

OpenStudy (lgbasallote):

\[\Large \int _1^{3^{37}} \frac{\pi \sin (\pi \ln x)}{x} dx\]

OpenStudy (lgbasallote):

?

OpenStudy (yash2651995):

e^37

OpenStudy (yash2651995):

lower limit corresponding to would be e^(1/pi)

OpenStudy (anonymous):

lower limit will be pi ln(1)=0

OpenStudy (yash2651995):

aacha ok... and upper be 37 pi sin t dt

OpenStudy (anonymous):

yes upper limit will be ..........37 pi

OpenStudy (yash2651995):

its 2 isnt it?

OpenStudy (anonymous):

thats right

OpenStudy (anonymous):

u did it well

OpenStudy (yash2651995):

you did it.. i cant believe i couldnt do such easy question in exam.. phew i have to keep a cool head.. -.- anyways thanks =)

OpenStudy (anonymous):

welcome

OpenStudy (yash2651995):

here another one.. tell me if i am right or wrong if f: D->R such that f(x) =sqrt(sin(cosx)) + ln(-2cos^2(x) + 3 cos x -1) then x1 to x2 integrate [cos(x) - 1/2] =? where x1 and x2 belongs to D and [.] is greatest integral part... my answer comes out to be "0"